OpenStudy (anonymous):

3x+4+6x-5+2x-8 what does x equal???

OpenStudy (saifoo.khan):

that expression must be equal to zero?

OpenStudy (anonymous):

Huh???

OpenStudy (saifoo.khan):

3x+4+6x-5+2x-8 = 0 <-- like that?

OpenStudy (anonymous):

no 3x+4+6x-5+2x-8=180

OpenStudy (saifoo.khan):

3x+4+6x-5+2x-8=180 Add like terms, 11x -9 = 180 11x = 180 + 9 11x = 189 x = 189/11

OpenStudy (anonymous):

the answer doesn't make sense /:

OpenStudy (saifoo.khan):

Why is that? check your equation, is it correct or not. :p

OpenStudy (anonymous):

yes, when you divide 189 by 11 it comes out un even

OpenStudy (anonymous):

the answer doesn't make sense /:

OpenStudy (anonymous):

the answer doesn't make sense /: